# MUTE-SLAM: Real-Time Neural SLAM with Multiple Tri-Plane Hash Representations

### v1

@@ 4.2 Evaluation of Mapping and Tracking
4.3 Performance Analysis
Table 6: Run-time and memory comparison on Replica [23], ScanNet [6], and Apartment [35] scenes. ESLAM [12] and Co-SLAM [28] were evaluated using their default settings from the open-source code. MUTE-SLAM demonstrates faster performance in middle and large scale scenes compared to ESLAM [12], and shows a slight edge in speed over Co-SLAM [28] on the large scale Apartment scene [35]. MUTE-SLAM also maintains a reasonable speed in the small scale synthetic ’room0’ scene from Replica [23].
@@ 4.3 Performance Analysis
Replica [23] ESLAM [12] 0.18 6.85M
Co-SLAM [28] 0.12 0.26M
@@
ScanNet [6] ESLAM [12] 0.56 17.8M
Co-SLAM [28] 0.19 1.59M
@@
Apartment [35] ESLAM [12] 2.40 22.1M
Co-SLAM [28] 0.23 1.59M
@@
We conducted a comparative analysis of the speed and memory consumption between our proposed MUTE-SLAM and the state-of-the-art methods ESLAM [12] and Co-SLAM [28]. Evaluations were performed on diverse scales of scenes: the small-scale ’room0’ from Replica [23], the mid-scale ’$0$’ from ScanNet [6], and the large-scale Apartment scene from NICE-SLAM [35]. Our metrics included average frame processing time (FPT) and the model’s parameter count. As shown in Tab. 6, MUTE-SLAM not only operates faster in large scale scenes but also requires less memory compared to ESLAM [12]. Notably, in large-scale scenarios like Apartment [35], MUTE-SLAM achieves even speed advantages over Co-SLAM [28]. Moreover, the FPT and memory usage of MUTE-SLAM remain relatively stable across scene sizes, a benefit attributable to our scene representation design.
4.4 Ablations
Table 7: Quantitative results of ablation study. Our full model achieves the best tracking and overall mapping performance.
Method Replica [23] ScanNet [6]
@@ 4.4 Ablations
4.4.1 Multi-map representation.
We conducted ablation experiments on the Replica [23] and ScanNet [6] datasets to evaluate the impact of various components of our design. The quantitative results of this study are detailed in Tab. 7. Representing the scene with one map, our findings indicate that the multi-map representation improves tracking performance. This enhancement in tracking accuracy, in turn, leads to higher quality in the reconstruction process. This improvement can be attributed to our strategy for allocating submaps. By extending the boundaries of each submap over a defined length, the corresponding hash tables are able to attain larger sizes, which contributes to better overall system performance.
4.4.2 Tri-plane hash-encoding.
To assess the effectiveness of tri-plane hash-encoding, we conducted an experiment where we replaced the tri-plane in each hash-encoding with a grid, while simultaneously tripling the maximum hash table size $N_{max}$. This adjustment marginally increases the overall capacity of the hash tables compared to the tri-plane approach. Tab. 7 shows that tri-plane hash-encoding achieves superior tracking results, a higher completion ratio, and better-rendered depth images. Although grid hash-encoding excels in terms of accuracy and completion, it leads to artifacts in the reconstructed mesh due to hash collisions, which in turn affects tracking accuracy. As illustrated in Fig. 5, our qualitative comparison demonstrates that our proposed tri-plane hash-encoding effectively reduces aliasing and preserves scene details more accurately.
4.4.3 Global bundle adjustment.
As illustrated in Tab. 7, the lack of global bundle adjustment leads to higher ATE errors and relatively low reconstruction performance. As global bundle adjustment corrects drifting poses and refines scene representation, it plays a critical role in ensuring robustness and global consistency in our method.
5 Conclusion

### v3

@@ IV-B4 Evaluation on Apartment [23]
IV-C Performance Analysis
TABLE VI: Run-time and memory comparison on Replica [35], ScanNet [36], and Apartment [23] scenes.
@@ IV-C Performance Analysis
Replica ESLAM 0.18 6.85M
Co-SLAM 0.12 0.26M
@@
ScanNet ESLAM 0.56 17.8M
Co-SLAM 0.19 1.59M
@@
Apartment ESLAM 2.40 22.1M
Co-SLAM 0.23 1.59M
@@
We conducted a comparative analysis of the speed and memory consumption between our proposed MUTE-SLAM and the state-of-the-art methods ESLAM [28] and Co-SLAM [24]. Evaluations were performed on diverse scales of scenes: the small-scale ’room0’ from Replica [35], the mid-scale ’$0$’ from ScanNet [36], and the large-scale Apartment scene from NICE-SLAM [23]. Our metrics included average frame processing time (FPT) and the model’s parameter count. As shown in table VI, MUTE-SLAM not only operates faster in large scale scenes but also requires less memory compared to ESLAM [28]. Notably, in large-scale scenarios like Apartment [23], MUTE-SLAM achieves even speed advantages over Co-SLAM [24]. Moreover, the FPT and memory usage of MUTE-SLAM remain relatively stable across scene sizes, a benefit attributable to our scene representation design. Although Co-SLAM’s [24] coordinate hash encoding reduces runtime and memory usage, its smoothing effect hinders detailed scene reconstruction. Therefore, coordinate hash encoding is preferable when fine-grained reconstructions are unnecessary, whereas our tri-plane hash offers a better balance of overhead and performance for detailed needs.
IV-D Ablations
TABLE VII: Quantitative results of ablation study.
Method Replica ScanNet
@@ IV-D Ablations
IV-D1 Multi-map representation
We conducted ablation experiments on the Replica [35] and ScanNet [36] datasets to evaluate the impact of various components of our design. The quantitative results of this study are detailed in table VII. Representing the scene with one map, our findings indicate that the multi-map representation improves tracking performance. This enhancement in tracking accuracy, in turn, leads to higher quality in the reconstruction process. This improvement can be attributed to our strategy for allocating submaps. By extending the boundaries of each submap over a defined length, the corresponding hash tables are able to attain larger sizes, which contributes to better overall system performance. To further prove that ill-set scene boundaries harm performances of mapping and tracking, we conducted a experiment on the Replica [35] ’room0’ sequence for all baselines in the need of pre-set scene boundaries as in table VIII. To simulate the scenario where the camera exits the pre-set boundary, we modify the boundary to a cube with a side length of 10m for all baselines (except for Vox-Fusion due to its implementation) and report the ’origin / ill-set’ results. Note that NICE-SLAM fails on this setting. The results indicate significant degradation for all baselines, meanwhile our multi-map representation will not encounter such a problem.
TABLE VIII: Impact of ill-set scene boundary.
Method Depth L1 $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$ Cp. Ratio $\uparrow$ ATE $\downarrow$
NICE-SLAM N/A N/A N/A N/A N/A
Vox-Fusion (64*0.2m) 1.12 / 3.45 1.21 / 1.48 1.35 / 1.61 93.85 / 90.42 1.2 / 6.7
ESLAM 0.95 / 49.74 1.04 / 1.11 1.03 / 33.97 99.71 / 72.04 0.67 / 9.3
Co-SLAM 1.47 / 52.09 1.04 / 1.24 1.06 / 31.47 99.45 / 72.74 0.63 / 0.64
IV-D2 Tri-plane hash-encoding
To assess the effectiveness of tri-plane hash-encoding, we conducted an experiment where we replaced the tri-plane in each hash-encoding with a grid, while simultaneously tripling the maximum hash table size $N_{max}$. This adjustment marginally increases the overall capacity of the hash tables compared to the tri-plane approach. table VII shows that tri-plane hash-encoding achieves superior tracking results, a higher completion ratio, and better-rendered depth images. Although grid hash-encoding excels in terms of accuracy and completion, it leads to artifacts in the reconstructed mesh due to hash collisions, which in turn affects tracking accuracy. As illustrated in fig. 5, our qualitative comparison demonstrates that our proposed tri-plane hash-encoding effectively reduces aliasing and preserves scene details more accurately.
IV-D3 Global bundle adjustment
As illustrated in table VII, the lack of global bundle adjustment leads to higher ATE errors and relatively low reconstruction performance. As global bundle adjustment corrects drifting poses and refines scene representation, it plays a critical role in ensuring robustness and global consistency in our method.
V Conclusion